OpenStudy (auw7):

Did I do this right? A treasure chest has a volume of x^3 + 9x^2 + 26x + 24 cm^3 and the height is x + 4 cm. Find the polynomial that would represent the area of the bottom of the treasure chest? Explain your reasoning. My steps: (x^3+9x^2+26x+24)/(x+4) x^2+5x+6 The polynomial that would represent the area of the bottom of the chest is (x+3)(x+2), because I factored the expression.

OpenStudy (auw7):

Will give a medal if anyone wants one lol

OpenStudy (anonymous):

I searched it and this is what I saw: the area of the bottom is A = (x^3 + 9x^2 + 26x + 24)/(x + 4) A = 6 + 5x + x^2

OpenStudy (benlindquist):

i want medal

OpenStudy (auw7):

@BraceFacee Thank you for your answer. And @benlindquist you'll have to answer the question of if you think it is correct if you want a medal lol!

OpenStudy (auw7):

I want to know if the last part is right or not.

OpenStudy (benlindquist):

yes

OpenStudy (auw7):

Are you sure or are you just saying that to get a medal?

OpenStudy (auw7):

Is my reasoning correct?

OpenStudy (benlindquist):

its correct 100% sure

OpenStudy (auw7):

k. I'll submit test in few minutes. Thanks